During the last few years, Jana Industries has been too constrained by the high cost of capital to make many capital investments. Recently, though, capital costs have been declining, and the company has decided to look seriously at a major expansion program that had been proposed by the marketing department. Assume that you are an assistant to Leigh Jones, the financial vice-president. Your first task is to estimate Jana's cost of capital. Jones has provided you with the following data, which she believes may be relevant to your task:

Information:

1. The firm's tax rate is 40 percent.

2. The current price of Harry Davis’ 12 percent coupon, semiannual payment, noncallable bonds with 15 years remaining to maturity is $1,153.72. Jana's does not use short-term interest-bearing debt on a permanent basis. New bonds would be privately placed with no flotation cost.

3. The current price of the firm's 10 percent, $100 par value, quarterly dividend, perpetual preferred stock is $116.95. Jana would incur flotation costs equal to 5% per share on a new issue.

4. Jana’s common stock is currently selling at $50 per share. Its last dividend (d0) was $3.12, and dividends are expected to grow at a constant rate of 5.8 percent in the foreseeable future. Jana’s beta is 1.2; the yield on t-bonds is 5.6 percent; and the market risk premium is estimated to be 6 percent. For the bond-yield-plus-risk-premium approach, the firm uses a 3.2 percentage point risk premium. 5. Harry Davis’ target capital structure is 30 percent long-term debt, 10 percent preferred stock, and 60 percent common equity.

5. Jana's target capital structure is 30% long-term debt, 10% preferred stock, and 60% common equity.

Problems:

A. What is the firms cost of preferred stock?

B. Jana's preferred stock is riskier to investors than its debt, yet the preferred stock's yield to investors is lower than the yield to maturity on the debt. Does this suggest that you have made a mistake? (hint think about taxes)

C. What is the estimated cost of equity using the discounted cash flow approach
Answer #1

New Microso cel Workshect Microsoft Excel HomeInsertPage LayoutFormulas Data Review View s Cut Σ AutoSum Copy B 1 프 田, Cr-소,

cost of preferred stock preferred dividend/net proceeds 10/111.95 8.93%
preferred stock 10
net proceeds 116.95-(100*5%) 111.95
before tax cost of debt = Using rate function in MS excel rate(nper,pmt,pv,fv,type) nper = 15*2 = 30 pmt = 1000*12%*1/2 = 60 pv =1153.72 fv =1000 type =0 RATE(30,-60,1153.72,-1000,0) 5.00%
annual tax before cost of debt 5*2 10%
after tax cost of debt before tax cost of debt*(1-tax rate) 10*(1-.4) 6
Yes cost of preferred stock is lower than before tax cost of bonds but we consider after tax cost of debt due to interest tax shield due to which cost of bond or debt reduce to. So after tax cost of debt is less than cost of preferred stock which is used as component cost to calculate WACC
cost of common stock with no floatation cost (expected dividend/market price)+growth rate (3.30/50)+5.8% 12.40%
expected dividend current dividend*(1+growth rate) 3.12*1.058 3.30096
market price 50
growth rate 5.80%
cost of common stock with floatation cost (expected dividend/market price)+growth rate (3.30/47.5)+5.8% 12.75%
expected dividend current dividend*(1+growth rate) 3.12*1.058 3.30096
net proceed from new issue 50*(1-.05) 47.5
growth rate 5.80%